OpenStudy (anonymous):

I really don't understand my lesson on lines and slope. (Edit: Can somebody answer the question at the very bottom?)

OpenStudy (anonymous):

Could somebody lead me through one problem and then check a second one that I try by myself?

OpenStudy (anonymous):

The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5). y = −2x + 13 y = negative 1 over 2x + 7 y = 1 over 2x + 3 y = − 2x − 3

OpenStudy (anonymous):

I would especially appreciate a step by-step explanation of what to do.

OpenStudy (anonymous):

A parallel line has the same slope Plug into the equation y-y1 = m(x-x1)

OpenStudy (anonymous):

slope = -2 point (4,5) y - 5 = -2(x-4) y-5 = -2x+8 y = -2x + 13 in standard form the line would be 2x+y = 13

OpenStudy (anonymous):

Slope intercept form is the y = -2x + 13

OpenStudy (anonymous):

ok give me a minute to comprehend that

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} y = &{\color{red}{ -2}}x − 2\\ &\ \uparrow \\ &slope \end{array}\)

OpenStudy (anonymous):

OK...if you want to do another problem, I'll watch what you do.

OpenStudy (anonymous):

^ thank you. @jdoe0001 I don't even know which bits represent what in the equation.

OpenStudy (anonymous):

And ok, @DSS . Just one second, thanks

OpenStudy (anonymous):

The -2 is the y-intercept...means the line crosses the y-axis at (0,-2)

OpenStudy (jdoe0001):

y = (whatever value here) x + (who knows some other value) ^ ^ slope y-intercept

OpenStudy (anonymous):

This is the picture of y = -2x-2

OpenStudy (jdoe0001):

so, you'd want to solve for "y" to get it to look like so first, so you can spot them easily

OpenStudy (anonymous):

and is the x the x intercept? or is the x intercept ot used?

OpenStudy (anonymous):

*not

OpenStudy (jdoe0001):

y-intercept

OpenStudy (anonymous):

the x that you didn't mark is the y intercept?

OpenStudy (jdoe0001):

no, the "x" is part of the equation, but is NEITHER the slope or the y-intercept, the "x" you can say is just along for the ride =)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

What are the rules about parallel and perpendicular lines? Like doesn't the perpendicular have to flip and turn the numerator negative?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

ok And then parallel stays the same for slope?

OpenStudy (anonymous):

oh ok wow

OpenStudy (anonymous):

the resources I'm provided with always astound me

OpenStudy (jdoe0001):

as you can see, parallel lines have the same slope perpendicular ones, have "negative reciprocal" slopes

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

I think I get it? I'll do my other question, could one of y'all tell me whether i get it right or not?

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

Thank you. It might take a minute.

OpenStudy (anonymous):

The bit that you labeled as Y intercept has a fraction in this problem. Why is that?

OpenStudy (anonymous):

The equation of line WX is y = −2x − 5. Write an equation of a line perpendicular to line WX in slope-intercept form that contains point (−1, −2). y = 1 over 2x + 3 over 2 y = negative 1 over 2x + 3 over 2 y = 1 over 2x − 3 over 2 y = − 1 over 2x − 3 over 2

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well... you know the slope of a perpendicular line to y = −2x − 5, right?

OpenStudy (anonymous):

Youd turn -2 into -2/1 and then flip it

OpenStudy (anonymous):

but then what bit would be negative?

OpenStudy (jdoe0001):

well... -2 into a +2 and flip it, yes 1/2

OpenStudy (anonymous):

oh ok. So you just change whatever the sign was before... so the 2 is positive

OpenStudy (jdoe0001):

pretty much, yes NEGATIVE RECIPROCAL a NEGATIVE OF a -a RECIPROCAL of that 1/-a

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the slope of the perpendicular line is 1/2?

OpenStudy (anonymous):

but wait, I thought it was the 1 that should change signs?

OpenStudy (jdoe0001):

y = −2x − 5 ^ slope -2 negative -(-2) = +2 reciprocal 1/2

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

So how do I know if the y intercept is positive or negative?

OpenStudy (jdoe0001):

now to find the equation of that line, now that we know its slope, well \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -1}}\quad ,&{\color{blue}{ -2}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{1}{2} \\ \quad \\ y-y_1={\color{green} m}(x-x_1)\Leftarrow\textit{point-slope form, plug in the values}\\ \quad \\ y-{\color{blue}{ -2}}={\color{green}{ \cfrac{1}{2}}}(x-{\color{red}{ -1}})\)

OpenStudy (jdoe0001):

then, distribute, simplify and solve for "y"

OpenStudy (anonymous):

Can i turn y--2 into y+2 ?

OpenStudy (anonymous):

Or will that somehow distort the answer?

OpenStudy (jdoe0001):

nope. you ARE expected to multiply them, yes, so y--2 or y-(-2) = y+2, that's correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got y+2 =1/2x-1/2 thus far

OpenStudy (jdoe0001):

hmm x-(-1) = x+1

OpenStudy (anonymous):

I'm confused

OpenStudy (jdoe0001):

\(\bf y-{\color{blue}{ -2}}={\color{green}{ \cfrac{1}{2}}}(x-{\color{red}{ -1}})\implies y+2= \cfrac{1}{2}(x+1)\)

OpenStudy (jdoe0001):

\(\bf y-({\color{blue}{ -2}})={\color{green}{ \cfrac{1}{2}}}(x-({\color{red}{ -1}}))\implies y+2= \cfrac{1}{2}(x+1)\)

OpenStudy (anonymous):

oh no i wrote it down wrong...

OpenStudy (anonymous):

y= 1/2 x - 1 1/2?

OpenStudy (anonymous):

Which would turn to 3/2

OpenStudy (jdoe0001):

yeap y= 1/2 x - 1 1/2\(\bf \checkmark\)

OpenStudy (anonymous):

ok awesome

OpenStudy (anonymous):

so that would be d?

OpenStudy (anonymous):

oh wait no

OpenStudy (jdoe0001):

slope is positive, so would be "c"

OpenStudy (anonymous):

c

OpenStudy (anonymous):

lol yeah

OpenStudy (anonymous):

ok great

OpenStudy (anonymous):

Thank you so much!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

And the first question The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5). was -----> y = −2x + 13 right?

OpenStudy (theeric):

Yep! :)